Please answer question 4.5
ran- options are available. For each of these three options, another three options are presented. If a person calls the 800 number for assistance, sam- mple how many total options are possible? t the 4.5. A bin contains six parts. Two of the parts are defective and four are acceptable. If three of the six parts are selected from the bin, how large is the sample space? Which counting rule did you use, and why? For this sample spac three sampled parts is defective? 4,7), e, what is the probability that exactly one of the
Answer #1

3 of 6 parts are selected. Here, order is not important

Thus, counting rule that we will use is combinations

Sample space size= 6C3 = 6! / (3!*3!) = 20

P(defective) = 2/6 = 1/3

P(acceptable) = 4/6 = 2/3

Using Binomial theorem,

P(one defective of 3) = 3C1*P(defective)1*P(acceptable)2

= 3 * (1/3) * (2/3)2

= 0.444